### v1

# Kobayashi hyperbolicity of complete linear systems on abelian varieties

Abstract.
We provide a geometric condition ensuring that a very general element of a complete linear system on an abelian variety is Kobayashi hyperbolic. Some related conjectures are also given.
1. Introduction
It is well-known that a complex abelian variety $A$ is not Kobayashi hyperbolic, as it admits non-constant entire curves, that is, non-constant holomorphic maps ${\mathbb{C}}\to A$. Besides, by the so-called Bloch’s Theorem, an entire curve $f\colon{\mathbb{C}}\to A$ is subject to a very strong restriction: the Zariski clousure $\overline{f({\mathbb{C}})}$ is a translation of an abelian subvariety of $A$. This result has a long history. It was stated in 1926 by Bloch [Bl] with an incomplete proof. A complete argument was given only around the end of the 70’s, independently and at the same time, by several people: Ochiai [Oc] and Kawamata [Ka], Smyth [Sm], Green-Griffiths [GG], and Wong [Wo]. All these proofs basically builds on the original approach of Bloch. McQuillan [McQ] later found a different, more arithmetic argument.
It is also known that the complement of an ample divisor on an abelian variety is Brody hyperbolic, [^1] As usual, one refers to a variety which admits no non-constant entire curves by saying that it is Brody hyperbolic. This condition is equivalent to the Kobayashi hyperbolicity for compact complex variety by Brody’s theorem [Br]. In general, Kobayashi hyperbolicity implies Brody’s one. that is, the image $f({\mathbb{C}})$ of a non-constant entire curve $f\colon{\mathbb{C}}\to A$ must intersect the support of any ample divisor on $A$. This was conjectured by Lang, and proved by Siu-Yeung [SY] (see also [Ya] for a stronger and more recent result). We refer the interested reader to [Ko, DR], and to [De, Ja] for surveys on Kobayashi hyperbolicity and related topics.
Very recently, the following result appeared:
Theorem 1.1 ([IMRY], Corollary 1.10).
Let $A$ be a complex abelian variety of dimension $g\geq 3$, and let $L$ be an ample line bundle on $A$. Then,
  1. 1)
    The linear system $|mL|$ is algebraically hyperbolic, if $m\geq g$.
  2. 2)
    If $|L|$ has no base divisors, then $m\geq g-1$ suffices.
A basepoint-free [^2] We recall that, on an abelian variety, $|mL|$ is always basepoint-free as soon as $m\geq 2$ (see, e.g., [BL, Proposition 4.1.5]). linear system is said to be algebraically (resp. Kobayashi) hyperbolic, if its very general element is algebraically (resp. Kobayashi) hyperbolic as a variety. Algebraic hyperbolicity is a notion introduced by Demailly in [De] as an algebraic analogue of Kobayashi hyperbolicity. The reader may find in §2 its definition and the (conjectural) relation with Kobayashi hyperbolicity. Although not noted in [IMRY], it is known that these two notions are equivalent for closed subvarieties of abelian varieties (see Remark 2.2 below). Theorem 1.1 confirms, in the case of abelian varieties and with actually a better bound, a general conjecture of Moraga and Yeong [MY] about the algebraic hyperbolicity of adjoint bundles (see Conjecture 2.3 below), which has been motivated by Fujita’s famous conjecture [F].
Let us recall that, thanks to the Decomposition Theorem (see [BL, §4.3], or [Ke2, §10.3]), the base locus of an ample line bundle $L$ on an abelian variety $A$ contains a divisor if and only if the polarized abelian variety $(A,L)$ contains a principally polarized factor $(B,\Theta)$, i.e., $(A,L)\simeq(A^{\prime},L^{\prime})\times(B,\Theta)$ as polarized abelian varieties. As observed in [IMRY], Item (1) of Theorem 1.1 is optimal: if
(1.1) $(A,L)\simeq(A^{\prime},L^{\prime})\times(E,\Theta),$
with $(E,\Theta)$ a principally polarized elliptic curve, then any element of $|mL|$ is not hyperbolic when $m\leq g-1$, and $g=\dim A\geq 2$ (see op.cit., Proof of Proposition 1.12). We point out that polarized abelian varieties as in (1.1) have been characterized by Nakamaye [Na], and, more recently, by Alvarado-Pareschi [AP2].
In this short note, we aim to show how taking into account the geometry of a polarized abelian variety $(A,L)$ leads to an improvement of Theorem 1.1, in such a way that (1.1) turns out to be the only “exceptional” case. Namely, we have:
Theorem A.
Let $A$ be a complex abelian variety of dimension $g\geq 3$, and let $L$ be an ample line bundle on $A$. Then, $|mL|$ is Kobayashi hyperbolic if
$m\geq g-1\,$,
unless $(A,L)$ is as in (1.1). In this case, one needs to take $m\geq g$.
The proof of Theorem A uses a criterion in [IMRY], which guarantees the algebraic hyperbolicity of a basepoint-free linear system on AA, [^3] We state it directly for abelian varieties as Theorem 2.4, but it holds true more generally for varieties with nef tangent bundle. This is used in [IMRY] to get the algebraic hyperbolicity of linear systems in several other cases. and recent results of Pareschi [Pa1, Pa2] and Alvarado-Pareschi [AP1, AP2] about certain generation properties of coherent sheaves on abelian varieties. In particular, we get
Special Case B.
Let $(A,\Theta)$ be an indecomposable principally polarized abelian variety of dimension $g\geq 3$. Then, the linear system $|m\Theta|$ is Kobayashi hyperbolic if $m\geq g-1$.
Here, indecomposable means that $(A,\Theta)$ cannot be written as the product of two lower-dimensional p.p.a.v.’s $(A_{i},\Theta_{i})$, which, accordingly to the above-mentioned Decomposition Theorem, is equivalent to saying that the theta divisor $\Theta$ is irreducible.
Concerning the sensibly simpler case of abelian surfaces, we have the next result, which is included for completeness.
Proposition C.
Let $A$ be a complex abelian variety of dimension $g=2$, and let $L$ be an ample line bundle on $A$. Then, $|mL|$ is Kobayashi hyperbolic as soon as $m\geq 2$, and $|L|$ itself is Kobayashi hyperbolic, if $L$ is basepoint-free.
Given a (possibly primitive) polarized abelian variety $(A,L)$ of dimension $g\geq 1$, a constant $\beta(A,L)$ attached to it, which is called the basepoint-freeness threshold and has been introduced in [JP], recently proved to detect a lot of information about the positivity of $(A,L)$ (see [JP, Ca, Ji1], to name just a few). In [Ca], the author conjectured that $\beta(A,L)$ could be bounded from above in terms of the degree (with respect to $L$) of all non-zero abelian subvarieties of $A$ (see (5.1)). This would provide a higher syzygies analogue, for abelian varieties, of the generalized Fujita’s conjecture. Along with the hyperbolicity criterion of [IMRY] (Criterion 2.4 below) and [IMRY, Theorem 1.9], this suggests a conjecture about the hyperbolicity of a (possibly primitive) linear system $|L|$ on an abelian variety $A$:
Conjecture D.
Let $(A,L)$ be a polarized abelian variety of dimension $g\geq 2$, such that
(1.2) $(L^{\dim B}\cdot B)>((g-1)\dim B)^{\dim B}$
for all abelian subvarieties $\{0\}\neq B\subseteq A$. Then, $|L|$ is Kobayashi hyperbolic.
We note that part of the above-mentioned conjecture of [Ca] predicts that, under the assumption (1.2), $L$ is indeed basepoint-free (and actually very ample, when $g\geq 3$). We will show in §5 that Conjecture D is true in dimension $g\leq 3$, and “almost true” if $g=4$. Moreover, it holds in arbitrary dimension “up to a factor 22” (see Proposition 5.3). We refer the reader to §5 for more on this circle of ideas. Here we just collect the next result for a general member of the moduli space of polarized abelian varieties of a given type, which follows from the semi-continuity of the basepoint-freeness threshold and from certain known upper bounds:
Proposition E.
Let $(A,L)$ be general polarized abelian variety of dimension $g\geq 2$. If
$h^{0}(A,L)\geq\frac{(2g(g-1))^{g}}{g!}\,$,
then $|L|$ is Kobayashi hyperbolic.
This already improves [IMRY, Corollary 1.13], and Proposition 5.5 below gives an even better (but more involved) result. Moreover, the following question arises naturally:
Question F.
Let $(A,L)$ be a general polarized abelian variety of dimension $g\geq 2$, with
$h^{0}(A,L)>(g-1)^{g}\,.$
Is the linear system $|L|$ Kobayashi hyperbolic?
Acknowledgment. I would like to thank Beppe Pareschi for his interest and for answering my questions about the proof of Lemma 4.1, and Ariyan Javanpeykar for Remark 2.2, and for his kind guidance through the hyperbolicity literature. The author is a member of GNSAGA-INdAM.
2. Algebraic hyperbolicity
As mentioned in the Introduction, the next definition has been introduced by Demailly [De] (see also [JK]).
Definition 2.1.
A complex projective variety $X$ is said to be algebraically hyperbolic if there exist a real number $\varepsilon>0$ and an ample line bundle $L$ on $X$ such that, for every smooth projective curve $C$ and every non-constant morphism $f\colon C\to X$, one has that
$2g(C)-2\geq\varepsilon\,\mathrm{deg}_{C}(f^{*}L),$
where $g(C)$ is the genus of $C$.
In particular, an algebraically hyperbolic variety does not contain any rational and elliptic curves. Demailly [De] proved that a Kobayashi hyperbolic projective variety is algebraically hyperbolic, and conjectured that also the converse implication holds true. This is, in general, quite open.
Since we are particularly interested in subvarieties of abelian varieties, let us promptly observe the following.
Remark 2.2.
Let $Y$ be a closed subvariety of an abelian variety $A$. The following are equivalent:
  1. 1)
    $Y$ is Kobayashi hyperbolic;
  2. 2)
    $Y$ is Brody hyperbolic;
  3. 3)
    $Y$ is algebraically hyperbolic;
  4. 4)
    $Y$ does not contain translates of positive-dimensional abelian subvarieties of $A$.
Indeed, (1) and (2) are equivalent thanks to Brody’s theorem [Br]. Item (1) implies (3) by Demailly’s theorem [De, Theorem 2.1] (see also the comment below [JK, Theorem 1.2]). The implication (3) $\Rightarrow$ (4) is, for instance, implicit in [JK]. Indeed, [JK, Corollary 4.5] says that an algebraically hyperbolic variety is groupless, that is, it does not admit any non-trivial map from an algebraic group. So, in particular, one gets (4). Finally, (4) implies (2) thanks to the Bloch’s theorem mentioned in the Introduction. The equivalence between (1) and (4) was also proved by Green in [Gr].
Quite naturally, an ample and basepoint-free complete linear system $|L|$ on $X$ is said to be algebraically hyperbolic, if a very general element of $|L|$ is algebraically hyperbolic. A similar definition may be given with the Kobayashi hyperbolicity. Moraga and Yeong conjectured what follows.
Conjecture 2.3 ([MY], Conjecture 1.1).
Let $L$ be an ample line bundle on a smooth projective variety $X$ of dimension $n\geq 2$. Then, the linear system $|K_{X}+(3n+1)L|$ is algebraically hyperbolic.
They have been motivated by Fujita’s conjecture [F], and by both classical and more recent results on hyperbolicity of (very) general hypersurfaces in projective spaces (see op.cit., the references therein, and [DR]). They proved their conjecture for smooth projective toric varieties. More recently, [IMRY] produced an useful criterion to detect the algebraic hyperbolicity of basepoint-free ample complete linear systems on a variety $X$ with nef tangent bundle (and possibly mild singularities). We state it here directly when $X=A$ is an abelian variety, and refer the reader to [IMRY, Theorem 4.7] for the more general statement. Its proof builds on a nowadays standard deformation type argument, recently simplified and further developed by Coskun and Riedl in [CR1, CR2].
Criterion 2.4 ([IMRY]).
Let $A$ be a complex abelian variety of dimension $g\geq 3$. Let $N$ and $L$ be line bundles on $A$, with $N$ globally generated and $L$ ample. If there exists a rational number $\delta>0$ such that
  1. i)
    $M_{N}{\langle}\delta L{\rangle}$ is nef, and
  2. ii)
    $N-\delta(g-2)L$ is ample,
then $|N|$ is an algebraically hyperbolic linear system on $A$.
Let us recall that $M_{N}$ is the kernel bundle of $N$, that is, the kernel of the evaluation morphism of global sections of NN
$0\to M_{N}\to H^{0}(A,N)\otimes{\mathcal{O}}_{A}\to N\to 0\,$,
and $M_{N}{\langle}\delta L{\rangle}$ denotes the $\mathbb{Q}$}-twisted bundle in the sense of Lazarsfeld [LaI, §6.2A].
We plain to apply the above Criterion 2.4 and Remark 2.2 to get Theorem A. With this in mind, the notions and results of the next section are useful. Before moving on to it, let us consider the easy case of hyperbolicity of linear systems on abelian surfaces.
Proof of Proposition C.
Let $A$ be an abelian surface, and $L$ be an ample and globally generated line bundle on $A$. Take $D\in|L|$ general, so that $D$ is a smooth projective irreducible curve. By taking the holomorphic Euler chracteristic of its defining short exact sequence
$0\to{\mathcal{O}}_{A}(-D)\to{\mathcal{O}}_{A}\to{\mathcal{O}}_{D}\to 0\,$,
we get
$0=\chi(A,{\mathcal{O}}_{A})$ $=\chi(D,{\mathcal{O}}_{D})+\chi(A,{\mathcal{O}}_{A}(-D))$
$=1-g(D)+h^{2}(A,{\mathcal{O}}_{A}(-D))\,$,
where $g(D)$ is the genus of $D$. Therefore, we have
$g(D)=1+h^{0}(A,{\mathcal{O}}_{A}(D))>1\,$,
as $D$ is an ample (hence effective) divisor on $A$. This means that $D$ is a hyperbolic curve.
If $L$ is not globally generated, we may apply the above reasoning to $|mL|$, which is basepoint-free when $m\geq 2$. This concludes the proof of Proposition C. ∎
3. Coherent sheaves on abelian varieties: regularity and generation properties
Let $(A,L)$ be a polarized abelian variety of dimension $g\geq 1$. Given a coherent sheaf $\mathcal{F}$ on $A$, we denote by
$V^{i}(A,\mathcal{F}):=\{\alpha\in\mathrm{Pic}^{0}A\ |\ h^{i}(A,\mathcal{F}\otimes\alpha)\neq 0\}$
the $i$-th cohomological support loci. These are Zariski closed subsets of $\mathrm{Pic}^{0}A$ by upper-semicontinuity. The sheaf $\mathcal{F}$ is said to be GV (resp. M-regular) if
$\mathrm{codim}_{\mathrm{Pic}^{0}A}V^{i}(A,\mathcal{F})\geq i\quad(\mathrm{resp.}>i)$
for all $i>0$.
Such notions turn out to be quite useful in the study of linear systems on abelian varieties (see [PP1]), and have been extended to the $\mathbb{Q}$}-twisted setting by Jiang-Pareschi [JP] as follows: given a $\mathbb{Q}$}-twisted sheaf $\mathcal{F}{\langle}rL{\rangle}$ on $A$, with $r=\frac{a}{b}\in{\mathbb{Q}}$ and $b>0$, we say that $\mathcal{F}{\langle}rL{\rangle}$ is GV (resp. M-regular) if so is $\mu_{b}^{*}(\mathcal{F})\otimes L^{ab}$, where
$\mu_{b}\colon A\to A$
is the multiplication-by-$b$ isogeny of $A$. The definition makes sense because $\mu_{b}^{*}L$ is algebraically equivalent to $L^{b^{2}}$, and hence $L^{ab}$ plays the rôle of $\mu_{b}^{*}(\frac{a}{b}L)$”.
Theorems of Debarre [De] and Pareschi-Popa [PP3] say that an $M$-regular sheaf is ample, and a $GV$ sheaf is nef, and it is clear that these implications remain valid in the $\mathbb{Q}$}-twisted setting (see, e.g., [IMRY, Lemma 3.7]). Note also that the converse implications are not true, unless we are dealing with line bundles. As observed in [IMRY, Remark 3.10], if $m\geq 2$, one has that the $\mathbb{Q}$}-twisted sheaf
(3.1) $M_{mL}{\langle}\frac{m}{m-1}L{\rangle}$
is $GV$, and hence it is nef. This is a consequence of the computations performed in [JP, §8], involving the Fourier-Mukai-Poincaré transform. However, (3.1) may not be $M$-regular, and actually it is so precisely when $L$ has no base divisors. Indeed, by [JP, Proposition 8.1] (see also [It2, Proposition 4.1]), one has that (3.1) is $M$-regular if and only if
$\mathcal{I}_{p}{\langle}\frac{1}{m}mL{\rangle}=\mathcal{I}_{p}\otimes L$
is $M$-regular, where $\mathcal{I}_{p}$ is the ideal sheaf of a closed point $p\in A$, and this is equivalent to ask that $L$ has no base divisors (see [PP2, Example 3.10(3)]).
Let us now explain, for the sake of completeness and self-containedness, how [IMRY] obtained Theorem 1.1 from the above results, by applying their Criterion 2.4 with $N=mL$.
Proof of Theorem 1.1.
Fix an integer $m\geq 2$. As we explained above,
(3.2) $M_{mL}{\langle}\frac{m}{m-1}L{\rangle}$
is nef, and it is ample if $L$ has no base divisors. Let us assume we are in the latter case, and take a rational number $\delta<\frac{m}{m-1}$ such that $M_{mL}{\langle}\delta L{\rangle}$ is still ample. So, we only need to check (ii) of the Criterion 2.4. Since
(3.3) $m-\delta(g-2)>m-\frac{m}{m-1}(g-2),$
one has that $mL-\delta(g-2)L=(m-\delta(g-2))L$ is ample as soon as the right-hand side of (3.3) is $\geq 0$, and this happens if and only if $m\geq g-1$.
If our line bundle $L$ has some base divisor, then we still have that (3.2) is nef, and we may take $\delta=\frac{m}{m-1}$. In this case the Criterion 2.4 applies as soon as $(m-\delta(g-2))L$ is ample, that is, as soon as
$m-\frac{m}{m-1}(g-2)>0\,$,
i.e., if $m\geq g$. ∎
Remark 3.1.
More generally, but with essentially the same argument, [IMRY, Theorem 1.9] proves that for a polarized abelian variety $(A,L)$ of dimension $g\geq 3$ and such that
(3.4) $\mathcal{I}_{p}{\langle}\frac{1}{g-1}L{\rangle}\quad\mathrm{is\,\emph{M-}regular,}$
one has that $|L|$ is very ample and hyperbolic. Indeed, (3.4) gives, in particular, that $L$ is globally generated, and one still has in this case, again by [JP], that (3.4) is equivalent to the $M$-regularity of $M_{L}{\langle}\frac{1}{g-2}L{\rangle}$. Then, the proof proceeds in the same way as before.
As observed (see (1.1)), Theorem 1.1(1) is optimal. However, we note that the above proof leaves room for improvement, as it is used that the $\mathbb{Q}$}-twisted sheaf (3.2) is $M$-regular (and hence ample), while what is really needed is just its ampleness. In this regard, we recall the following
Definition 3.2 ([Pa2]).
Let $A$ be an abelian variety of dimension $g\geq 1$, and $Z$ be an irreducible subvariety of $\mathrm{Pic}^{0}A$. A coherent sheaf $\mathcal{F}$ on $A$ is said to be generated by $Z$ if the sum of twisted evaluation maps
$\mathrm{ev}_{U}\colon\bigoplus_{\alpha\in U}H^{0}(A,\mathcal{F}\otimes\alpha)\otimes\alpha^{\vee}\to\mathcal{F}$
is surjective for all non-empty open subset $U\subseteq\mathrm{Pic}^{0}A$ meeting $Z$.
A generated sheaf is nef, as it can be written as the quotient of a finite direct sum of numerically trivial line bundles (see [Pa2, Remark 2.1.4]). A theorem of Debarre [De] can be stated, using this language, by saying that an $M$-regular sheaf on $A$ is generated by the full $\mathrm{Pic}^{0}A$ (this is what is commonly called being continuously globally generated). Finally, let us note that a globally generated sheaf is generated by $\{\hat{0}\}\subseteq\mathrm{Pic}^{0}A$, but the converse does not hold in general.
In [Pa2], Pareschi gave an ampleness criterion for generated sheaves. [^4] We state below a rather special case, which suffices to our aims. Let us say that an irreducible subvariety $Z\subseteq\mathrm{Pic}^{0}A$ spans $\mathrm{Pic}^{0}A$ if there exists an integer $N\geq 1$ such that
$\underbrace{Z+\ldots+Z}_{N}=\mathrm{Pic}^{0}A\,.$
Theorem 3.3 ([Pa2], Theorem A).
Let $\mathcal{F}$ be a coherent sheaf on $A$. If $\mathcal{F}$ is generated by an irreducible subvariety spanning $\mathrm{Pic}^{0}A$, then $\mathcal{F}$ is ample.
So, going back to our previous notations, we observed that $M_{mL}{\langle}\frac{m}{m-1}L{\rangle}$ is always $GV$ when $m\geq 2$, and now the problem is to understand when it is generated by a spanning subvariety of $\mathrm{Pic}^{0}A$.
4. Proof of Theorem A
Let us start be considering a polarized abelian variety $(A,L)$ of dimension $g\geq 1$, for the moment. We need to introduce some more notations. In [Pa1, AP1], the authors noted (and used) that the $\mathbb{Q}$}-twisted sheaf ${\mathcal{O}}_{A}{\langle}rL{\rangle}$, with $r\in{\mathbb{Q}}$, cohomologically behaves as a certain “usual” vector bundle, denoted by $E_{A,rL}$. This is a simple semi-homogeneous vector bundle on $A$ such that
$\frac{\det E_{A,rL}}{\mathrm{rk\,}E_{A,rL}}=rL$
holds in the Néron-Severi group $\mathrm{NS}(A)_{{\mathbb{Q}}}$ of $A$. Its existence was proved by Mukai [Mu1, §7] (see also [AP1, §1.5]). It has the following crucial property [AP1, Proposition 2.1.1]:
(4.1) $\mu_{b}^{*}E_{A,\frac{a}{b}L}\simeq({L^{\prime}}^{ab})^{\oplus\mathrm{rk\,}E_{A,\frac{a}{b}L}}\,$,
where $L^{\prime}$ is a line bundle on $A$, algebraically equivalent to $L$. This fact says us that, for a $\mathbb{Q}$}-twisted sheaf $\mathcal{F}{\langle}\frac{a}{b}L{\rangle}$, considering the sheaf $\mu_{b}^{*}(\mathcal{F})\otimes L^{ab}$ as in §3 is not the only option to deal with an “integral” sheaf, instead of a $\mathbb{Q}$}-twisted one. Another, cohomologically equivalent, possibility is to look at $\mathcal{F}\otimes E_{A,\frac{a}{b}L}$, because
$\mu_{b}^{*}(\mathcal{F}^{\,\oplus\mathrm{rk\,}(E_{A,rL})}\otimes E_{A,rL})=(\mu_{b}^{*}\mathcal{F}\otimes{L^{\prime}}^{ab})^{\oplus\mathrm{rk\,}E_{A,rL}}\,.$
We refer the reader to [AP1], where the theory coming from this observation is fully developed.
With this terminology in our hands, we may now recall a recent result of Alvarado-Pareschi [AP2], which has its roots in [Pa1].
Lemma 4.1 ([AP2]).
Let $(A,\Theta)$ be an indecomposable p.p.a.v. of dimension $g\geq 2$. Then, for any integer $m\geq 2$, the sheaf
$M_{m\Theta}\otimes E_{A,\frac{m}{m-1}\Theta}$
is generated by an ample prime divisor $D$ of $\mathrm{Pic}^{0}A$.
Since this Lemma is proved in [AP2, Lemma 3.2.1] with $m=2$ (where $E_{A,2\Theta}=2\Theta$), and the details are left to the reader when $m\geq 3$ (see [AP2, §5.2]), we will include a sketch of its proof in the Appendix below.
Corollary 4.2.
Under the same assumptions, the $\mathbb{Q}$}-twisted sheaf $M_{m\Theta}{\langle}\frac{m}{m-1}\Theta{\rangle}$ is ample.
Proof.
By the above Lemma 4.1 and Theorem 3.3, $M_{m\Theta}\otimes E_{A,\frac{m}{m-1}\Theta}$ is an ample bundle, if the generating ample prime divisor $D$ spans $\mathrm{Pic}^{0}A$. But taking cohomology in
$0\to{\mathcal{O}}_{\mathrm{Pic}^{0}A}(-D)\to{\mathcal{O}}_{\mathrm{Pic}^{0}A}\to{\mathcal{O}}_{D}\to 0$
gives
$0=H^{g-1}({\mathcal{O}}_{\mathrm{Pic}^{0}A}(D))=H^{1}({\mathcal{O}}_{\mathrm{Pic}^{0}A}(-D))\to H^{1}({\mathcal{O}}_{\mathrm{Pic}^{0}A})\to H^{1}({\mathcal{O}}_{D})\,$,
that is, $h^{1}({\mathcal{O}}_{D})\geq H^{1}({\mathcal{O}}_{\mathrm{Pic}^{0}A})=g>g-1=\dim D$. Hence, $D$ cannot be a translation of an abelian subvariety of $\mathrm{Pic}^{0}A$. It follows thereby (see, e.g., [Ue, p. 109]) that, for dimensional reasons, $D$ spans $\mathrm{Pic}^{0}A$.
Now, thanks to (4.1), we have
(4.2) $\mu_{m-1}^{*}(M_{m\Theta}\otimes E_{A,\frac{m}{m-1}\Theta})\simeq\mu_{m-1}^{*}(M_{m\Theta})\otimes({{\Theta}^{\prime}}^{\,m(m-1)})^{\oplus\mathrm{rk\,}E_{A,\frac{m}{m-1}\Theta}}\,.$
Therefore, by standard properties on ampleness of vector bundles, we get that
$\mu_{m-1}^{*}(M_{m\Theta})\otimes{\Theta}^{\,m(m-1)}=\mu_{m-1}^{*}(M_{m\Theta}{\langle}\frac{m}{m-1}\Theta{\rangle})$
is ample. Since $\mu_{m-1}\colon A\to A$ is finite and surjective, one has that $M_{m\Theta}{\langle}\frac{m}{m-1}\Theta{\rangle}$ is ample as well (see [LaII, Lemma 6.2.8]). ∎
Proof of the special case B.
It follows at once from the above Corollary 4.2 and the Criterion 2.4, as in the proof of Theorem 1.1. Namely, we take an indecomposable p.p.a.v. $(A,\Theta)$ of dimension $g\geq 3$. Thanks to Corollary 4.2, there exists a rational number $0<\delta<\frac{m}{m-1}$, where $m\geq 2$, such that $M_{m\Theta}{\langle}\delta\Theta{\rangle}$ is ample. Since, as soon as $m\geq g-1$, one has $m-\delta(g-2)>m-\frac{m}{m-1}(g-2)\geq 0$, we get that $(m-\delta(g-2))\Theta$ is ample. Therefore, we may apply the Criterion 2.4 – and Remark 2.2 – to $|m\Theta|$, if $m\geq g-1$. ∎
We are finally ready to give the
Proof of Theorem A.
Let $(A,L)$ be a polarized abelian variety of dimension $g\geq 3$, that is not isomorphic, as a polarized abelian variety, to a product of a principally polarized elliptic curve with another polarized abelian variety as in (1.1). By the Decomposition Theorem ([BL, §4.3], or [Ke2, §10.3]), we have that
$(A,L)\simeq(B,N)\times(A_{1},\Theta_{1})\times\ldots\times(A_{k},\Theta_{k})\,$,
where $N$ is an ample line bundle without base divisors on an abelian variety $B$, and the $(A_{i},\Theta_{i})$’s are indecomposable p.p.a.v.’s of dimension $\geq 2$.
Thanks to the discussion below (3.1), one has that, if $m\geq 2$, $M_{mN}{\langle}\frac{m}{m-1}N{\rangle}$ is $M$-regular, and hence ample, on $B$. On the other hand, thanks to Corollary 4.2, $M_{m\Theta_{i}}{\langle}\frac{m}{m-1}\Theta_{i}{\rangle}$ is ample on $A_{i}$ for all $i=1,\ldots,k$. Let $0<\delta<\frac{m}{m-1}$ be a rational number such that all these $\mathbb{Q}$}-twisted bundles are still ample. Then, it follows easily (see, e.g., [IMRY, Proposition 3.4]) that
$M_{m(N\,\boxtimes\,\Theta_{1}\,\boxtimes\,\ldots\,\boxtimes\,\Theta_{k})}{\langle}\delta(N\boxtimes\Theta_{1}\boxtimes\ldots\boxtimes\Theta_{k}){\rangle}\simeq M_{mL}{\langle}\delta L{\rangle}$
is a nef $\mathbb{Q}$}-twisted bundle. Since $m-\delta(g-2)>m-\frac{m}{m-1}(g-2)\geq 0$, where the second inequality holds true as we are assuming $m\geq g-1$, we get that
$mL-\delta(g-2)L$
is ample. Hence, it only rests to apply the Criterion 2.4 of [IMRY], and Remark 2.2. ∎
5. The basepoint-freeness threshold and hyperbolicity of linear systems
Let $(A,L)$ be a polarized abelian variety of dimension $g\geq 1$, and $\mathcal{I}_{p}$ be the ideal sheaf of a closed point $p\in A$. Then, the basepoint-freeness threshold of $(A,L)$ is
$\beta(A,L):=\mathrm{Inf}\{r\in{\mathbb{Q}}_{>0}\ |\ \mathcal{I}_{p}{\langle}rL{\rangle}\,\textrm{is \emph{GV}}\}\,.$
It is known that $L$ is basepoint-free if and only if $\beta(A,L)<1$, and moreover, if $\beta(A,L)<r_{0}$, then $\mathcal{I}_{p}{\langle}r_{0}L{\rangle}$ is $M$-regular (see [JP, Theorem 5.2]). Hence, we have
Remark 5.1.
If $\beta(A,L)<\frac{1}{g-1}$, with $g\geq 2$, then $|L|$ is hyperbolic. Indeed, this follows from Proposition C when $g=2$, and from [IMRY, Theorem 1.9] (see also Remark 3.1 above) when $g\geq 3$.
On the other hand, it is expected by [Ca] that the following geometric upper bound holds
(5.1) $\beta(A,L)\leq\mathrm{Max}_{B}\,\frac{\dim B}{\sqrt[\dim B]{((L^{\dim B}\cdot B)}}\,$,
where $B$ varies among all non-zero abelian subvarieties of AA. [^5] Note that we really have a Max and not only a Sup here. This is because, given $A$, there are, up to isomorphism, only finitely many abelian subvarieties in $A$ (see [LOZ]). So we reduce to deal with two isomorphic $d$-dimensional abelian subvarieties $B$ and $B^{\prime}$ of $A$ such that $(L^{d}\cdot B)\neq(L^{d}\cdot B^{\prime})$. The $B^{\prime}$’s that are relevant to the computing of the Sup in (5.1) are just those with $(L^{d}\cdot B)>(L^{d}\cdot B^{\prime})$. Therefore, once $B$ is fixed, we only have a finite number of possibilities for the quotient appearing in the Sup. This has been verified for abelian surfaces [Ca], for abelian $3$-folds [It1], “almost” for abelian $4$-folds [Ji2, Theorem 1.6], and in arbitrary dimension “up to a factor 22[Ji1, Theorem 1.4]. Namely,
Theorem 5.2.
  1. 1)[Ji2]
    If $g=4$, one has
    $\beta(A,L)\leq\mathrm{Max}\{\frac{4.31}{\sqrt[4]{((L^{4})}},\frac{\dim B}{\sqrt[\dim B]{((L^{\dim B}\cdot B)}}\ |\ \{0\}\neq B\subsetneq A\}\,;$
  2. 2)[Ji1]
    When $g$ is arbitrary,
    (5.2) $\beta(A,L)\leq\mathrm{Max}_{B}\,\frac{2\,\dim B}{\sqrt[\dim B]{((L^{\dim B}\cdot B)}}\,.$
Therefore, by requiring that the right-hand side of (5.2) is $\frac{1}{g-1}$, one directly obtains:
Proposition 5.3.
Let $(A,L)$ be a polarized abelian variety of dimension $g\geq 2$. If
$(L^{\dim B}\cdot B)>(2\,(g-1)\dim B)^{\dim B}$
for all abelian subvarieties $\{0\}\neq B\subseteq A$, then $|L|$ is Kobayashi hyperbolic.
In the same vein, Conjecture D of the Introduction would follow at once from (5.1), and hence it is essentially verified when $g\leq 4$.
In [Ji2], Jiang also proved the following estimate for $\beta(A,L)$, when $A$ is simple:
Theorem 5.4 ([Ji2], Theorems 1.6, 1.8 and 1.9).
If $(A,L)$ is a simple polarized abelian variety of dimension $g\geq 4$, there exists an explicit real constant $\alpha_{g}>0$ such that
(5.3) $\beta(A,L)\leq\frac{2g-\alpha_{g}}{\sqrt[g]{((L^{g})}}\,.$
Even if, on a simple abelian variety, any basepoint-free complete linear system is hyperbolic by Remark 2.2, such bound is useful to get results about the hyperbolicity of linear systems on general polarized abelian varieties.
Proposition 5.5.
Let $(A,L)$ be a general polarized abelian variety of dimension $g\geq 4$ and type $(d_{1},\ldots,d_{g})$. If
$h^{0}(A,L)>\frac{((2g-\alpha_{g})(g-1))^{g}}{g!}\,$,
then $|L|$ is Kobayashi hyperbolic.
Proof.
By Theorem 5.4, if $(A_{0},L_{0})$ is a simple polarized abelian variety of dimension $g$ and type $(d_{1},\ldots,d_{g})$, then
$\beta(A_{0},L_{0})\leq\frac{2g-\alpha_{g}}{\sqrt[g]{(L_{0}^{g})}}=\frac{2g-\alpha_{g}}{\sqrt[g]{g!\,h^{0}(A,L)}}\leq r\,$,
where $r<\frac{1}{g-1}$ is a rational number whose existence is guaranteed by our assumption. It is not difficult to see that the basepoint-freeness threshold is upper semi-continuous (see, e.g., [It3, Theorem 3.1]), hence
$\beta(A,L)\leq r$
for any general polarized abelian variety $(A,L)$ of dimension $g$ and of type $(d_{1},\ldots,d_{g})$. Since $r<\frac{1}{g-1}$, we get the hyperbolicity of $|L|$ by Remark 5.1. ∎
In particular, one obtains at once Proposition E of the Introduction. Similar (and sometimes better) results can also be obtained from [Ji2, Theorem 3.1].
When $(A,L)$ is general in its moduli space, it is also expected (see [Ji2, p. 3]) that
$\beta(A,L)\sim\frac{1}{\sqrt[g]{h^{0}(A,L)}}\,.$
Therefore, the following question seems natural.
Question 5.6.
Let $(A,L)$ be a general polarized abelian variety of dimension $g\geq 2$, with
$h^{0}(A,L)>(g-1)^{g}\,.$
Is the linear system $|L|$ Kobayashi hyperbolic?
Appendix: Proof of Lemma 4.1
As said above, since the proof of Lemma 4.1 is given in [AP2, Lemma 3.2.1] just with $m=2$ (in which case $E_{A,2\Theta}=2\Theta$), and the details for $m\geq 3$ are left to the interested reader (see [AP2, §5.2]), we will give a sketch of their beautiful argument, for the reader’s benefit.
Lemma 5.7 ([AP2]).
Let $(A,\Theta)$ be an indecomposable p.p.a.v. of dimension $g\geq 2$. For any $m\geq 2$, the sheaf $M_{m\Theta}\otimes E_{A,\frac{m}{m-1}\Theta}$ is generated, in the sense of Definition 3.2, by a prime divisor in $\mathrm{Pic}^{0}A$, which is algebraically equivalent to $m^{2(g-1)}\widehat{\Theta}$, where $\widehat{\Theta}$ denotes the dual principal polarization on $\mathrm{Pic}^{0}A$.
Proof.
Consider the twisted evaluation map
$\mathrm{twe}\colon H^{0}(A,E_{A,\frac{m}{m-1}\Theta})\otimes{\mathcal{O}}_{A}(m\Theta)\to E_{A,\frac{m}{m-1}\Theta}(m\Theta)\simeq E_{A,(\frac{m}{m-1}+m)\Theta}=E_{A,\frac{m^{2}}{m-1}\Theta}\,$,
and denote by
$\Phi_{\mathcal{P}}\colon\mathrm{D}(A)\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}\mathrm{D}(\mathrm{Pic}^{0}A)$
the Fourier-Mukai-Poincaré equivalence [Mu2] (here $\mathcal{P}$ is the normalized Poincaré line bundle on $A\times\mathrm{Pic}^{0}A$, as usual). It follows from (4.1), thanks to cohomology and base change, that $\Phi_{\mathcal{P}}(E_{A,\frac{a}{b}\Theta})$ is a locally free sheaf on $\mathrm{Pic}^{0}A$, if $a,b>0$, and moreover the morphism
$\Phi_{\mathcal{P}}(\mathrm{twe})\colon H^{0}(A,E_{A,\frac{m}{m-1}\Theta})\otimes\Phi_{\mathcal{P}}({\mathcal{O}}_{A}(m\Theta))\to\Phi_{\mathcal{P}}(E_{A,\frac{m^{2}}{m-1}\Theta})$
of locally free sheaves on $\mathrm{Pic}^{0}A$ is fiber-wise the multiplication of global sections
(5.4) $\nu_{\alpha}\colon H^{0}(A,E_{A,\frac{m}{m-1}\Theta})\otimes H^{0}(A,{\mathcal{O}}_{A}(m\Theta)\otimes\alpha)\to H^{0}(A,E_{A,\frac{m^{2}}{m-1}\Theta}\otimes\alpha)\,$,
with $\alpha\in\mathrm{Pic}^{0}A$.
Let us assume (this can be harmlessly done, up to a translation) that the theta divisor $\Theta$ is symmetric. We now use the basepoint-freeness threshold (see §5). Since we deal with a principal polarization, it is not difficult to see that $\beta(m\Theta)=\frac{\beta(\Theta)}{m}=\frac{1}{m}$. Therefore, since $m\geq 2$, it follows from [JP, Proposition 8.1] that
$M_{m\Theta}{\langle}\frac{m}{m-1}\Theta{\rangle}$
is $GV$. Thanks to (4.2), this means that $M_{m\Theta}\otimes E_{A,\frac{m}{m-1}\Theta}$ is $GV$ too, which in turn implies that (5.4) are surjective for general $\alpha\in\mathrm{Pic}^{0}A$. Since the dimension of the source and of the target of (5.4) are equal (to $m^{2g}$, see [Mu1, Theorem 7.11(5)], or [AP1, (1.9)]), the locus of the $\alpha$’s such that $\nu_{\alpha}$ drops rank is a divisor. It coincides with the support of the divisor of zeroes of the determinant morphism $\det(\Phi_{\mathcal{P}}(\mathrm{twe}))$.
Let us denote this divisor of zeroes by $D$. The Néron-Severi class of $D$ is the one of $m^{2(g-1)}\widehat{\Theta}$, where $\widehat{\Theta}$ is the dual principal polarization on $\mathrm{Pic}^{0}A$. To do this computation, one uses the fact the first Chern class of the source and of the target of the morphism $\det(\Phi_{\mathcal{P}}(\mathrm{twe}))$ can be easily calculated using the fact (see, e.g., [AP1, Proposition 2.2.1]) that
(5.5) $\Phi_{\mathcal{P}}(E_{A,\frac{a}{b}\Theta})=E_{\mathrm{Pic}^{0}A,\frac{-b}{a}\widehat{\Theta}}$
when $a>0$, and moreover that (see [AP1, (1.9)])
$\mathrm{rk\,}(E_{\mathrm{Pic}^{0}A,\frac{-1}{m}\widehat{\Theta}})=m^{g}\quad\,\mathrm{and}\quad\mathrm{rk\,}(E_{\mathrm{Pic}^{0}A,\frac{-(m-1)}{m^{2}}\widehat{\Theta}})=m^{2g}\,.$
The next step is to show that $D$ is indeed a prime divisor. Having this in mind, let us consider the pullback morphism
[…]
(5.6) $E=\sum_{x\in A[m]}\Theta_{x}\,$,
where $\Theta_{x}:=\Theta-x$ and $A[m]$ is the subgroup of $m$-torsion points of $A$. Note that, since $\Theta$ is supposed to be irreducible, on the right-hand side we have a sum of distinct prime divisors. The equality (5.6) is a consequence of [Pa1, Theorem A], which generalizes a classical result of Kempf [Ke1, Theorem 3]. Indeed, the morphism $\varphi_{m\Theta}^{*}(\Phi_{\mathcal{P}}(\mathrm{twe}))$ of sheaves on $A$, fiber-wise is the multiplication map of global sections
$\nu_{p}\colon H^{0}(A,E_{A,\frac{m}{m-1}\Theta})\otimes H^{0}(A,t_{p}^{*}{\mathcal{O}}_{A}(m\Theta))\to H^{0}(A,E_{A,\frac{m}{m-1}\Theta}\otimes t_{p}^{*}{\mathcal{O}}_{A}(m\Theta))\,$,
with $p\in A$. Now, [Pa1, Theorem A] [^7] In [Pa1], $W_{b,a}$ denotes what here is called $E_{A,\frac{a}{b}\Theta}$. says that the dimension of the cokernel of $\nu_{p}$ is precisely the cardinality of
$A[m]\cap\Theta_{p}\,$,
which is the set of $m$-torsion points sitting in the translated theta divisor $\Theta_{p}$. Therefore, the map $\nu_{p}$ is not surjective if and only if there exists a point $x\in A[m]\cap\Theta_{p}$, which is equivalent to say that $p\in\Theta_{x}$ for a point $x\in A[m]$. This suffices to say that the support of $E$ equals the support of the right-hand side of (5.6). Since we already showed that the Néron-Severi class of $E$ is the same of the one of the right-hand sum of (5.6), we get the equality in (5.6).
Now, since $\varphi_{m\Theta}^{*}{\varphi_{m\Theta}}_{*}(\Theta)=\sum_{x\in A[m]}\Theta_{x}$ and since $\varphi_{m\Theta}^{*}D=E$, one has that, up to translation by an $m$-torsion point,
$D={\varphi_{m\Theta}}_{*}(\Theta)\,$,
and it is known that ${\varphi_{m\Theta}}_{*}(\Theta)$ is a prime divisor (see [Pa1, p. 1591]).
Finally, the proof of the generation of the $GV$ sheaf $M_{m\Theta}\otimes E_{A,\frac{m}{m-1}\Theta}$ by the prime divisor $D$ goes as in [AP2, §3.3, Proof of Lemma 3.2.1(2)], and it is based on [Pa2, Corollary 6.2.1] (see also [AP2, §3.4] for a self-contained proof of this result). ∎